OpenStudy (anonymous):

How do you write (3,-3) m=-3/8 in standard form?

OpenStudy (arthur_ser):

y=mx+c -3=(-3/8)*3 + c -3=-9/8 + c c=-3 + 9/8 c=-24/8 + 9/8 c=-15/8 -->y=-(3/8)x -(15/8)